OpenStudy (anonymous):

This one may be a bit tricky, but it's the last one for today: find the coefficient o f (x^4)(y^3) in the expansion of (2x+y)^7.

OpenStudy (anonymous):

\[{7! \over 4!(3!)}(2x)^4y^3=35(16)x^4y^3=560x^4y^3\]